OpenStudy (anonymous):

at t=0, a particle starts at the origin with a velocity of 6ft persecond and moves along the x-axis in such a way that at time t its acceleration is 12t^2 ft per second per second. through how many feet does the particle move during the first 2 seconds?

OpenStudy (anonymous):

S=UT+(AT^2)/2 ; S - DISTANCE , U- INITIAL VELOCITY , A- ACCELERATION ; = 6*T+(12T^2)/2 =6*T+6*T^2=6*T(1+T)=6*2(1+2)=36 FT